OpenStudy (anonymous):

what is (sqrt)m12 simplified?

OpenStudy (anonymous):

m a variable?

OpenStudy (anonymous):

is it m^12?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\sqrt{12m}\] \[2\sqrt{3m}\]

OpenStudy (mertsj):

\[2\sqrt{3m}\]

OpenStudy (mertsj):

m^6 then if it is m^12

OpenStudy (anonymous):

lol thanks for stealing my fire lol

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Who stole what :D

OpenStudy (mertsj):

\[\sqrt{m ^{12}}=m ^{6}\]

OpenStudy (anonymous):

well technically its eaquals m^12/2= m^6

OpenStudy (anonymous):

Yeah so, \[\sqrt{m ^{12}}\] is same as: \[\sqrt{x ^{6^{2}}}\] So take square roots, we end up with \[x ^{6}\]